The length of a bicycle pedal arm is 0.173 m, and a downward force of 116 N is applied to the pedal by the rider. What is the magnitude of the torque about the pedal arm's pivot when the arm is at an angle of (a)43.2°, (b) 90°, and (c) 180° with the vertical?
Answer #1

Torque:

T=rF sin\theta ---------1

a)

T= 0.173×116× sin 43.2 =13.73 Nm

b)

T=0.173×116× sin 90 =20.068 Nm

c)

T= 0.173×116× sin 180 = 0 Nm